# MHE under parametric uncertainty - Robust state estimation without informative data

### v1

@@ Abstract
In this paper, we study state estimation for general nonlinear systems with unknown parameters and persistent process and measurement noise. In particular, we are interested in stability properties of the state estimate in the absence of persistency of excitation (PE). With a simple academic example, we show that existing moving horizon estimation (MHE) approaches as well as classical adaptive observers can result in diverging state estimates in the absence of PE, even if the noise is small. We propose a novel MHE formulation involving a regularization based on a constant prior estimate of the unknown system parameters. Only assuming the existence of a stable estimator, we prove that the proposed MHE results in practically robustly stable state estimates even in the absence of PE. We discuss the relation of the proposed MHE formulation to state-of-the-art results from MHE, adaptive estimation, and functional estimation. The properties of the proposed MHE approach are illustrated with a numerical example of a car with unknown tire friction parameters.
@@ 1 Introduction
State estimation is of high importance for control, system monitoring, and fault detection in many applications such as power systems [1], battery management systems [2], manufacturing systems [3], or biomedical systems [4]. Obtaining an accurate estimate of the state in general requires access to a model which accurately describes the underlying system dynamics. Identifying the parameters of such a model exactly is, however, difficult, e.g., due to the nonlinear nature of the models. In addition, the parameters are often slowly varying, e.g., due to aging of components, and therefore practical applications require some form of online adaptation of the model parameters. This renders the simultaneous estimation of states and parameters of nonlinear systems from noisy output measurements a problem of high practical relevance, compare, e.g, [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, the resulting estimation problem is nonlinear and non-convex even in the simple case of linear systems, see, e.g., [16, 17], and thus challenging to solve. Furthermore, unique identification of model parameters requires sufficiently informative data, typically through a so-called persistency of excitation (PE) condition (cf. [18, 19, 20, 21]), even if noise-free state measurements are available. Reliable parameter estimation is not possible in the absence of PE or, more generally, if the excitation is small compared to the noise, i.e., the signal to noise ratio is small. For general nonlinear systems or over-parametrized models, ensuring PE is typically impossible. For linear systems, PE can be ensured using additional excitation [20], however, this requires active deterioration of the regulation performance. In this paper, we present an estimation method with guaranteed robust stability for general nonlinear systems with unknown parameters subject to process and measurement noise, without relying on a PE assumption.
@@
More recently, joint state and parameter estimation for nonlinear systems has been addressed using optimization-based approaches, see, e.g., [37, 38] for the case of linear systems. Moving horizon estimation (MHE) is a promising optimization-based technique for estimation in general nonlinear systems subject to process and measurement noise [39, 40, 41]. In particular, uniform [^1] Uniform in the control input. robust stability of MHE can be established with a sufficiently large horizon under a general detectability condition in the form of incremental input/output-to-state-stability ($\delta$-IOSS) [42, 40]. Compared to classical estimation approaches, MHE requires us to solve a finite-horizon optimization problem online, which lead to the development of efficient solution strategies [10, 43, 44, 45]. Joint state and parameter estimation in MHE is commonly addressed by introducing an augmented state containing both the original system states and the unknown parameters [46, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Such MHE approaches are employed in different applications reaching from autonomous driving and friction estimation [5, 6], estimation in vibrating structures [7, 8], chemical processes [9, 10, 11, 12], biological processes [13, 14], to induction motors [15]. Uniform detectability ($\delta$-IOSS) of the augmented state consisting of the system state and parameters in general requires a uniform PE condition  [47, 48]. Consequently, the (robust) stability properties of these MHE schemes deteriorate arbitrarily if the excitation becomes small during online operation. Additionally, applying an MHE for joint state and parameter estimation in the absence of PE can result in numerical issues. While this has been addressed by modifications of the prior weighting in the MHE objective, compare [48, 49], the resulting MHE schemes do not provide any stability guarantees in the absence of PE. Using recent results from functional estimation [50], stability of the state estimate can be ensured even if the augmented state is not detectable. However, this requires the use of a full information estimator (FIE), i.e., an optimization-based estimator with unbounded horizon. The design of finite horizon MHE approaches with robust stability properties in the absence of PE for general nonlinear systems with parametric uncertainty and process and measurement noise remains an open question.
@@
In this paper, we derive an MHE formulation for general nonlinear systems with parametric uncertainty subject to process and measurement noise which ensures reliable state estimation in the absence of PE, and consequently without unique identifiability of the parameters. With a simple academic example, we show that standard MHE schemes for joint state and parameter estimation can lead to parameter drift in the absence of PE and thus lack fundamental robust stability properties (Section 3). We also demonstrate that classical adaptive observers have the same limitation (Section 5.3). To overcome this, we propose a novel MHE design by replacing the standard prior weighting of the parameter estimate with a simple regularization (Section 4). We show practical robust stability of the resulting state estimation error under the proposed MHE approach. Robust stability implies that the estimation error is proportional to the process and measurement noise. The stability is practical with respect to (w.r.t.) a constant involving the bias of the parameter regularization, which can be made arbitrarily small by increasing the horizon. To establish robust stability we only require a weak detectability condition on the system: There exists an estimator satisfying the desired robust stability property. Notably, satisfaction of this assumption does not require detectability of the augmented system state and hence does not impose a restrictive PE condition. The presented theoretical analysis relies on recent techniques developed in the context of MHE [40] and generalizations thereof to functional estimation [50]. We discuss the relation of the proposed MHE approach to classical MHE schemes and existing prior weightings to handle lack of PE, connect the proposed ”regularization” to classical tools from adaptive control and estimation, and discuss the relations to functional estimation (Section 5). Finally, we present a numerical example of a car with unknown tire friction parameters and in the absence of PE (Section 6).
@@
Let the non-negative real numbers be denoted by $\mathbb{R}_{\geq 0}$, the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a$ for some $a\in\mathbb{R}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for some $a,b\in\mathbb{R}$ with $a\leq b$ by $\mathbb{I}_{[a,b]}$. Let $\|x\|$ denote the Euclidean norm of the vector $x\in\mathbb{R}^{n}$. A positive definite (semi-definite) matrix $P=P^{\top}$ is denoted as $P\succ 0$ ($P\succeq 0$). The quadratic norm with respect to a positive definite matrix $Q=Q^{\top}$ is denoted by $\|x\|_{Q}^{2}=x^{\top}Qx$, and the minimal and maximal eigenvalues of $Q$ are denoted by $\lambda_{\min}(Q)$ and $\lambda_{\max}(Q)$, respectively. The maximum generalized eigenvalue of positive definite matrices $A=A^{\top}$ and $B=B^{\top}$ is denoted as $\lambda_{\max}(A,B)$, i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The $i$-th element of a vector $x\in\mathbb{R}^{n}$ is denoted as $[x]_{i}$ for $i\in\mathbb{I}_{[1,n]}$.
@@ 2.1 Problem Formulation
where $t\in\mathbb{I}_{\geq 0}$ is the discrete time step, $x_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$ the system state, $u_{t}\in\mathbb{U}=\mathbb{R}^{m}$ the control input, $w_{t}\in\mathbb{R}^{q}$ the process and measurement noise, $\theta\in\mathbb{R}^{n_{\theta}}$ a constant but unknown parameter, and $y_{t}\in\mathbb{Y}=\mathbb{R}^{p}$ the measurement output. Note that $w_{t}$ appears in the dynamics (1a) and measurement model (1b) and hence can also model separate process disturbances and measurement noise.
@@
with $\mathbb{X}\subseteq\mathbb{R}^{n_{x}}$, $\mathbb{W}\subseteq\mathbb{R}^{q}$, and $\Theta\subseteq\mathbb{R}^{n_{\theta}}$. Note that this is not restrictive, since we can always choose $\mathbb{X}=\mathbb{R}^{n_{x}}$, $\mathbb{W}=\mathbb{R}^{q}$, and $\Theta=\mathbb{R}^{n_{\theta}}$ in case no additional system knowledge is available. Constraints of the form (2) can be used to include additional physical prior knowledge, e.g., non-negativity of certain states, or bounds on noise and parameters, into a state estimator, compare [39, Sec. 4.3.3]. In the following, we assume that the true trajectory of system (2.1) always satisfies the constraints (2).
Given estimates $\bar{x}_{0}$ and $\bar{\theta}_{0}$ of the initial state $x_{0}$ and the unknown parameter $\theta$, respectively, the objective is to obtain an estimate $\hat{x}_{t}$ of the current system state $x_{t}$ at each time step $t\in\mathbb{I}_{\geq 0}$. The considered problem is formalized by the following definition of a state estimator.
Definition 1 (State estimator, adapted from [51, Def. 2.2])
A state estimator is a sequence of functions $\Psi_{t}:\mathbb{X}\times\Theta\times\mathbb{W}^{t}\times\mathbb{Y}^{t}\times\rightarrow\mathbb{X}$ for $t\in\mathbb{I}_{\geq 0}$ to compute estimates of (1a) at each time step $t$ as
@@
In a practical application, the noise estimates $\bar{w}_{t}$ are typically chosen to be zero and the output measurement $\bar{y}_{j}$ equal to the noisy output measurement $y_{j}$ according to (1b). The aim is to design a state estimator according to Definition 1 for system (2.1) which satisfies the following stability notion.
@@
$\|x_{t}-\hat{x}_{t}\|\leq$ $\max\left\{C_{1}\lambda_{1}^{t}\left\|\begin{bmatrix}x_{0}\\ \theta\end{bmatrix}-\begin{bmatrix}\bar{x}_{0}\\ \bar{\theta}_{0}\end{bmatrix}\right\|\right.,$ (8)
@@
This stability notion is an extension of the one defined in [50, Def. 3] in the context of functional estimation, compare also [52, Def. 6] for classical input-to-state practical stability. In contrast to classical stability, Inequality (9) only provides a bound on the estimation error of the state $x_{t}$, but not the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$. Note that in case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, and (9) is satisfied with $\epsilon=0$, Definition 2 corresponds to a standard robust stability notion considered in MHE, compare [40, Def. 2]. To design an $\delta$-IOpS state estimator, we require the following assumption toward the dynamics of the considered system.
@@ 2.2 Detectability
In the following, we introduce an incremental input/output-to-output-stability ($\delta$-IOOS) Lyapunov function as introduced in [50], which characterizes existence of a stable estimator (Assumption 1).
Definition 3 ($\delta$-IOOS Lyapunov function [50, Def. 6])
A function $W_{\delta}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{\theta}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}_{\geq 0}$ is an exponential $\delta$-IOOS Lyapunov function if there exist $P_{1},\ P_{2,x},\ P_{2,\theta}\succ 0$ and $Q,\ R\succeq 0$ such that
$\|x-\tilde{x}\|_{P_{1}}^{2}\leq W_{\delta}(x,\tilde{x},\theta,\tilde{\theta})\leq\|x-\tilde{x}\|_{P_{2,x}}^{2}+\|\theta-\tilde{\theta}\|_{P_{2,\theta}}^{2},$ (13a)
$W_{\delta}\left(x^{+},\tilde{x}^{+},\theta,\tilde{\theta}\right)\leq\eta W_{\delta}(x,\tilde{x},\theta,\tilde{\theta})$
@@
for all $\{x,u,w,y\}\ \in\ \mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{Y}$, $\{\tilde{x},u,\tilde{w},\tilde{y}\}\ \in\ \mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{Y}$, $\theta\in\Theta$, and $\tilde{\theta}\in\Theta$, where $x^{+}=f(x,u,w,\theta)$, $\tilde{x}^{+}=f(\tilde{x},u,\tilde{w},\tilde{\theta})$, $y=h(x,u,w,\theta)$, and $\tilde{y}=h(\tilde{x},u,\tilde{w},\tilde{\theta})$.
In the following proposition, we establish that Assumption 1 is a necessary and sufficient condition for the existence of a $\delta$-IOOS Lyapunov function according to Definition 3.
@@
Assumption 1 holds if and only if system (2.1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3.
@@
The considered problem is a special case of functional estimation studied in [50] with system state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ and corresponding functional output $z_{t}=\phi(x_{t},\theta)=x_{t}$ to be estimated. Resulting from [50, Cor. 8], system (2.1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3 if and only if the system is exponentially $\delta$-IOOS according to [50, Def. 4]. Additionally, it was shown in [50] that exponential $\delta$-IOOS [^2] Both [50, Prop. 5] and [50, Cor. 10] consider a more general asymptotic form of $\delta$-IOOS, but the proof naturally extends to the case of exponential $\delta$-IOOS. is both necessary (cf. [50, Prop. 5]) and sufficient (cf. [50, Cor. 10]) for Assumption 1 to be satisfied. Consequently, Assumption 1 is necessary and sufficient for the existence of an exponential $\delta$-IOOS Lyapunov function.
Existing stability results for MHE instead use a $\delta$-IOSS Lyapunov function, which would require $W_{\delta}$ to be lower bounded in the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$, cf. [51][40]. In [47], it was shown that this detectability condition also holds for the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ if $x_{t}$ is detectable and an additional a PE condition holds. In contrast, the detectability condition considered here does not depend on a PE condition, and thus holds uniformly for any sequence of inputs.
Note that we restrict our analysis to an exponential-like version of $\delta$-IOOS here. In [50, Thm. 9] a more general asymptotic version of $\delta$-IOOS was used to establish stability of a FIE approach for functional estimation. Notably, also without parametric uncertainty, establishing stability for an MHE scheme with finite horizon requires a more restrictive (exponential) condition, compare [41, Sec. 5.5.3] and [40, Rem. 4].

### v2

# MHE under parametric uncertainty — Robust state estimation without informative data

@@ Abstract
In this paper, we study joint state and parameter estimation for general nonlinear systems with uncertain parameters and persistent process and measurement noise. In particular, we are interested in stability properties of the resulting state estimate in the absence of persistency of excitation (PE). With a simple academic example, we show that existing moving horizon estimation (MHE) approaches for joint state and parameter estimation as well as classical adaptive observers can result in diverging state estimates in the absence of PE, even if the noise is small. We propose an MHE formulation involving a regularization based on a constant prior estimate of the unknown system parameters. Only assuming the existence of a stable state estimator, we prove that the proposed MHE approach results in practically robustly stable state estimates irrespective of PE. We discuss the relation of the proposed MHE formulation to state-of-the-art results from MHE and adaptive estimation. The properties of the proposed MHE approach are illustrated with a numerical example of a car with unknown tire friction parameters.
@@ 1 Introduction
State estimation is of high importance for control, system monitoring, and fault detection in many applications such as power systems [1], battery management systems [2], manufacturing systems [3], or biomedical systems [4]. Obtaining an accurate estimate of the state in general requires access to a model which accurately describes the underlying system dynamics. Identifying the parameters of such a model exactly is, however, difficult, e.g., due to the nonlinear nature of the models. In addition, the parameters are often slowly varying, e.g., due to aging of components, and therefore practical applications require some form of online adaptation of the model parameters. This renders simultaneous estimation of states and parameters of nonlinear systems from noisy output measurements a problem of high practical relevance, compare, e.g, [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, the resulting estimation problem is nonlinear and non-convex even in the simple case of linear systems, see, e.g., [17, 18], and thus challenging to solve. Furthermore, unique identification of model parameters requires sufficiently informative data, typically through a so-called persistency of excitation (PE) condition (cf. [19, 20, 21, 22]), even if noise-free state measurements are available. Reliable parameter estimation is not possible in the absence of PE or, more generally, if the excitation is small compared to the noise, i.e., the signal to noise ratio is small. For general nonlinear systems or over-parametrized models, ensuring PE is typically impossible. For linear systems, PE can be ensured using additional excitation [21], however, this requires active deterioration of the regulation performance. In this paper, we present an estimation method with guaranteed robust stability for general nonlinear systems with unknown parameters subject to process and measurement noise, without relying on a PE assumption.
@@
More recently, joint state and parameter estimation for nonlinear systems has been addressed using optimization-based approaches, see, e.g., [38, 39] for the case of linear systems. Moving horizon estimation (MHE) is a promising optimization-based technique for estimation in general nonlinear systems subject to process and measurement noise [40, 41, 42]. In case of perfectly known model parameters, uniform [^1] Uniform in the control input. robust stability of MHE for state estimation can be established with a sufficiently large horizon under a general detectability condition in the form of incremental input/output-to-state-stability ($\delta$-IOSS) [43, 41]. In this setting, it has been shown recently that $\delta$-IOSS is a necessary and sufficient condition for the existence of a robustly stable state estimator [44]. In addition, MHE has been shown to perform well even in situations where standard approaches like an extended Kalman filter (EKF) lead to non-physical state estimates [45]. Compared to classical estimation approaches, MHE relies on increased computational power to solve a finite-horizon optimization problem online. Consequently, this lead to the development of efficient solution strategies [10, 46, 47, 48]. Joint state and parameter estimation in MHE is commonly addressed by introducing an augmented state containing both the original system states and the unknown parameters [49, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Such MHE approaches are employed in different applications ranging from autonomous driving and friction estimation [5, 6], estimation in vibrating structures [7, 8], chemical processes [9, 10, 11, 12], biomedical systems [13, 14, 15], to induction motors [16]. Uniform detectability ($\delta$-IOSS) of the augmented state consisting of the system state and parameters in general requires a uniform PE condition [50, 51], which can often not be ensured in practice [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Consequently, the (robust) stability properties of these MHE schemes deteriorate arbitrarily if the excitation becomes small during online operation. Additionally, applying an MHE for joint state and parameter estimation in the absence of PE can result in numerical issues. While this has been addressed by modifications of the prior weighting in the MHE objective, compare [50, 52], the resulting MHE schemes do not provide any stability guarantees in the absence of PE. An MHE for joint estimation of states and parameters of a neural network system model, which generally suffers from lack of PE due to over-parametrization, involving a regularization on the parameter estimate was recently proposed in [53]. However, theoretical properties of the resulting estimates were not established. Using recent results from functional estimation [54], stability of the state estimate can be ensured even if the augmented state is not detectable. However, this requires the use of a full information estimator (FIE), i.e., an optimization-based estimator with unbounded horizon. The design of finite horizon MHE approaches with robust stability properties in the absence of PE for general nonlinear systems with parametric uncertainty and process and measurement noise remains an open question. Note that an MHE approach for joint state and parameter estimation, which ensures bounded state estimation errors in the absence of PE, was proposed concurrently in [51]. While the proposed work uses a necessary and sufficient detectability condition, the robust state detectability condition required in [51, Ass. 1] is generally more restrictive, compare also our discussion in Section 2.2 below.
@@
In this paper, we derive an MHE formulation for general nonlinear systems with parametric uncertainty subject to process and measurement noise which ensures reliable state estimation in the absence of PE, and consequently without unique identifiability of the parameters. With a simple academic example, we show that standard MHE schemes for joint state and parameter estimation can lead to parameter drift in the absence of PE and thus lack fundamental robust stability properties (Section 3). We also demonstrate that classical adaptive observers have the same limitation (Section 5.3). To overcome this, we propose to replace the standard prior weighting on the parameter estimate within the MHE formulation by a simple regularization involving a (constant) prior parameter estimate (Section 4). We show practical robust stability of the resulting state estimation error under the proposed MHE approach. Robust stability implies that the estimation error is proportional to the process and measurement noise. The stability is practical with respect to (w.r.t.) a constant involving the bias of the parameter regularization, which can be made arbitrarily small by increasing the horizon. To establish robust stability, we only require a weak detectability condition on the system: There exists an estimator satisfying the desired robust stability property. Hence, the proposed MHE approach is applicable whenever any robustly stable estimator for the considered system exists. Notably, satisfaction of this assumption does not require detectability of the augmented system state and hence does not impose a restrictive PE condition. The presented theoretical analysis relies on recent techniques developed in the context of MHE [41] and generalizations thereof to functional estimation [54]. We discuss the relation of the proposed MHE approach to classical MHE schemes and existing prior weightings to handle lack of PE, and connect the proposed “regularization” to classical tools from adaptive control and estimation (Section 5). Finally, we present a numerical example of a car with unknown tire friction parameters, with insufficient excitation over certain time periods (Section 6).
@@
Let the non-negative real numbers be denoted by $\mathbb{R}_{\geq 0}$, the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a$ for some $a\in\mathbb{R}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for some $a,b\in\mathbb{R}$ with $a\leq b$ by $\mathbb{I}_{[a,b]}$. Let $\|x\|$ denote the Euclidean norm of the vector $x\in\mathbb{R}^{n}$. A positive definite (semi-definite) matrix $P=P^{\top}$ is denoted as $P\succ 0$ ($P\succeq 0$). The quadratic norm with respect to a positive definite matrix $Q=Q^{\top}$ is denoted by $\|x\|_{Q}^{2}=x^{\top}Qx$, and the minimal and maximal eigenvalues of $Q$ are denoted by $\lambda_{\min}(Q)$ and $\lambda_{\max}(Q)$, respectively. The maximum generalized eigenvalue of positive definite matrices $A=A^{\top}$ and $B=B^{\top}$ is denoted as $\lambda_{\max}(A,B)$, i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The $i$-th element of a vector $x\in\mathbb{R}^{n}$ is denoted as $x]_{i}$ for $i\in\mathbb{I}_{[1,n]}$.
@@ 2.1 Problem Formulation
where $t\in\mathbb{I}_{\geq 0}$ is the discrete time step, $x_{t}\in\mathbb{R}^{n_{\mathrm{x}}}$ the system state, $u_{t}\in\mathbb{U}=\mathbb{R}^{n_{\mathrm{u}}}$ the control input, $w_{t}\in\mathbb{R}^{n_{\mathrm{w}}}$ the process and measurement noise, $\theta\in\mathbb{R}^{n_{\theta}}$ a constant but unknown parameter, and $y_{t}\in\mathbb{Y}=\mathbb{R}^{n_{\mathrm{y}}}$ the measurement output. Note that $w_{t}$ appears in the dynamics (1a) and measurement model (1b) and hence can also model separate process disturbances and measurement noise.
@@
with $\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$, $\mathbb{W}\subseteq\mathbb{R}^{n_{\mathrm{w}}}$, and $\Theta\subseteq\mathbb{R}^{n_{\theta}}$. Note that this is not restrictive, since we can always choose $\mathbb{X}=\mathbb{R}^{n_{\mathrm{x}}}$, $\mathbb{W}=\mathbb{R}^{n_{\mathrm{w}}}$, and $\Theta=\mathbb{R}^{n_{\theta}}$ in case no additional system knowledge is available. Constraints of the form (2) can be used to include additional physical prior knowledge, e.g., non-negativity of certain states, or bounds on noise and parameters, into a state estimator, compare [40, Sec. 4.3.3]. In the following, we assume that the true trajectory of system (1) always satisfies the constraints (2).
Given some available prior estimates $\bar{x}_{0}$ and $\bar{\theta}_{0}$ of the initial state $x_{0}$ and the unknown parameter $\theta$, respectively, the objective is to obtain an estimate $\hat{x}_{t}$ of the current system state $x_{t}$ at each time step $t\in\mathbb{I}_{\geq 0}$. The considered problem is formalized by the following definition of a state estimator.
Definition 1 (State estimator, adapted from [44, Def. 2.2])
A state estimator is a sequence of functions $\Psi_{t}:\mathbb{X}\times\Theta\times\mathbb{W}^{t}\times\mathbb{Y}^{t}\rightarrow\mathbb{X}$ for $t\in\mathbb{I}_{\geq 0}$ to compute estimates of (1a) at each time step $t$ as
@@
In a practical application, the noise estimates $\bar{w}_{t}$ are typically chosen to be zero and the output measurement $\bar{y}_{j}$ equal to the noisy output measurement $y_{j}$ according to (1b). The aim is to design a state estimator according to Definition 1 for system (1) which satisfies the following stability notion.
@@
$\|x_{t}-\hat{x}_{t}\|\leq$ $\max\left\{C_{1}\lambda_{1}^{t}\left\|\begin{bmatrix}x_{0}\\ \theta\end{bmatrix}-\begin{bmatrix}\bar{x}_{0}\\ \bar{\theta}_{0}\end{bmatrix}\right\|\right$.,
@@
This stability notion is an extension of the one defined in [54, Def. 3] in the context of functional estimation, compare also [55, Def. 6] for classical input-to-state practical stability. In contrast to classical stability, Inequality (4) only provides a bound on the estimation error of the state $x_{t}$, but not the augmented state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$. Note that in case the true parameter is known, i.e., $\bar{\theta}_{0}=\theta$, and (4) is satisfied with $\epsilon=0$, Definition 2 corresponds to a standard robust stability notion considered in MHE, compare [41, Def. 2]. To design an $\delta$-IOpS state estimator, we require the following assumption toward the dynamics of the considered system.
@@ 2.2 Detectability
In the following, we introduce an incremental input/output-to-output-stability ($\delta$-IOOS) Lyapunov function, which characterizes existence of a stable estimator (Assumption 1).
Definition 3 ($\delta$-IOOS Lyapunov function [54, Def. 6])
A function $W_{\delta}:\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\mathrm{x}}}\times\mathbb{R}^{n_{\theta}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}_{\geq 0}$ is an exponential $\delta$-IOOS Lyapunov function if there exist $P_{1},\ P_{2,\mathrm{x}},\ P_{2,\theta}\succ 0$, $Q,\ R\succeq 0$, and $\eta\in[0,1)$ such that
$\|x-\tilde{x}\|_{P_{1}}^{2}\leq W_{\delta}(x,\tilde{x},\theta,\tilde{\theta})\leq\|x-\tilde{x}\|_{P_{2,\mathrm{x}}}^{2}+\|\theta-\tilde{\theta}\|_{P_{2,\theta}}^{2},$ (5a)
[…]
@@
[…]
The following proposition shows that the existence of a $\delta$-IOOS Lyapunov function according to Definition 3 for the system (1) is a necessary and sufficient condition for Assumption 1 to be satisfied for the considered system.
@@
Assumption 1 holds if and only if system (1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3.
@@
The considered problem is a special case of functional estimation studied in [54] with system state $\begin{bmatrix}x_{t}^{\top}&\theta^{\top}\end{bmatrix}^{\top}$ and corresponding functional output $z_{t}=\phi(x_{t},\theta)=x_{t}$ to be estimated. Resulting from [54, Cor. 8], system (1) admits an exponential $\delta$-IOOS Lyapunov function according to Definition 3 if and only if the system is exponentially $\delta$-IOOS according to [54, Def. 4]. Additionally, it was shown in [54] that exponential $\delta$-IOOS [^2] Both [54, Prop. 5] and [54, Cor. 10] consider a more general asymptotic form of $\delta$-IOOS, but the proof naturally extends to the case of exponential $\delta$-IOOS. is both necessary (cf. [54, Prop. 5]) and sufficient (cf. [54, Cor. 10]) for Assumption 1 to be satisfied. Consequently, Assumption 1 is necessary and sufficient for the existence of an exponential $\delta$-IOOS Lyapunov function.
Note that we restrict our analysis to an exponential-like version of $\delta$-IOOS here. In [54, Thm. 9] a more general asymptotic version of $\delta$-IOOS was used to establish stability of a FIE approach for functional estimation. Notably, also without parametric uncertainty, establishing stability for an MHE scheme with finite horizon generally requires a more restrictive (exponential) condition, compare [42, Sec. 5.5.3] and [41, Rem. 4].
A related detectability condition was introduced concurrently in [51, Ass. 1]. Since the detectability condition invoked by Assumption 1 is necessary and sufficient, the condition in [51] cannot be less restrictive. In particular, [51, Ass. 1] considers a robust state detectability, which uses a $\delta$-IOSS Lyapunov function for the state and holds uniformly and robustly w.r.t. the parameters $\theta$. In contrast, Proposition 1 shows that the existence of a stable estimator under parametric uncertainty is equivalent to a $\delta$-IOOS Lyapunov function of the augmented state $(x,\theta)$. In the trivial case of a single possible model parameter $\Theta=\{\theta\}$, these two definitions coincide and reduce to a classical state detectability condition [44], but for the considered problem of uncertain model parameters they differ.